# VisualQuality-R1: Reasoning-Induced Image Quality Assessment via Reinforcement Learning to Rank

### v1

@@ 4.3 Ablation Studies
We vary the number of generated responses, $K$, while keeping all other settings fixed during GRPO. Table 3 shows that reducing $K$ from six (default) to four or five has only a marginal effect, offering a favorable trade-off between computational cost and accuracy.
@@
Table 4 shows that, within the same RL2R framework, our continuous fidelity reward generalizes better than the binary reward adopted in GRPO shao2024deepseekmath . Moreover, both reward variants consistently outperform the regression-based Q-Insight li2025qinsight , underscoring the effectiveness of our RL2R optimization.
Table 4: Comparison of different Thurstone model variants thurstone1927law in the GRPO of VisualQuality-R1.
@@ 4.4 Further Analysis
Figure 3: Prediction variability decreases during GRPO. We randomly select $20$ images from each of CLIVE ghadiyaram2015massive , KonIQ-10k hosu2020koniq , SRIQA-Bench chen2025toward , and AGIQA-3K li2023agiqa . At successive training steps, we generate multiple responses per image, compute the std of the predicted quality scores, and plot the average std across images. The uniformly downward trend confirms that VisualQuality-R1 becomes steadily more stable in assessing image quality as training progresses.
@@
We randomly sample $20$ images from each of CLIVE ghadiyaram2015massive , KonIQ-10k hosu2020koniq , SRIQA-Bench (super-resolution) chen2025toward , and AGIQA-3K (image generation) li2023agiqa , respectively. At successive training checkpoints, we generate multiple responses per image and compute the standard deviation (std) of the resulting $K$ quality scores. As illustrated in Fig. 3, the std falls steadily across all datasets, indicating that predictions of VisualQuality-R1 become progressively more stable and confident. This confirms that RL2R effectively reduces predictive uncertainty, improving the sensitivity of VisualQuality-R1 to subtle quality differences.
@@ 5 Conclusion and Discussion
Despite the generalization capabilities demonstrated by VisualQuality-R1, several limitations and promising research directions merit further discussion. First, as a specific case of test-time scaling, VisualQuality-R1 is slow, expensive, and memory-hungry; it may also compound early errors into confidently wrong predictions. It is thus desirable to incorporate sample-adaptive reasoning, rationale compression or distillation, and self-consistency sampling to make VisualQuality-R1 faster, cheaper, and more robust. Second, VisualQuality-R1 relies on a single, fixed text prompt for all images, regardless of the underlying distortion scenario or application context. Incorporating application-aware prompt adaptation, for example via learned prompt-tuning or dynamic template selection, could tailor VisualQuality-R1’s reasoning and scoring to specific image processing tasks, therefore improving its flexibility and accuracy. Third, VisualQuality-R1 is currently formulated as an NR-IQA model, focusing solely on distorted inputs without access to pristine counterparts. It is interesting to extend VisualQuality-R1 to a reference-based setting, which allows a (possibly corrupted) reference image wu2024comprehensive —potentially differing in resolution, color gamut, dynamic range, or bit depth—to serve as a flexible anchor for content fidelity. Last, we foresee adapting the proposed RL2R learning algorithm to other perceptual assessment tasks, including image aesthetics assessment tang2013content , human age estimation guo2009human , and perceptual similarity ranking fu2023dreamsim . Collectively, these promising directions aspire to foster more intelligent, transparent, and adaptable perceptual systems.

### v2

@@ 4.3 Ablation Studies
We vary the number of generated responses, $K$, while keeping all other settings fixed during GRPO. Table 3 shows that reducing $K$ from six (default) to four or five has only a marginal effect, offering a favorable trade‑off between computational cost and accuracy.
@@
Table 4 shows that, within the same RL2R framework, our continuous fidelity reward generalizes better than the binary reward adopted in GRPO shao2024deepseekmath . Moreover, both reward variants consistently outperform the regression‐based Q-Insight li2025qinsight , underscoring the effectiveness of our RL2R optimization.
Table 4: Comparison of different Thurstone model variants thurstone1927law in GRPO of VisualQuality-R1.
@@ 4.4 Further Analysis
Figure 3: Prediction variability decreases during GRPO. We randomly select 2020 images from each of CLIVE ghadiyaram2015massive , KonIQ-10k hosu2020koniq , SRIQA-Bench chen2025toward , and AGIQA-3K li2023agiqa . At successive training steps, we generate multiple responses per image, compute the std of the predicted quality scores, and plot the average std across images. The uniformly downward trend confirms that VisualQuality-R1 becomes steadily more stable in assessing image quality as training progresses.
@@
We randomly sample $20$ images from each of CLIVE ghadiyaram2015massive , KonIQ-10k hosu2020koniq , SRIQA-Bench (super-resolution) chen2025toward , and AGIQA-3K (image generation) li2023agiqa , respectively. At successive training checkpoints, we generate multiple responses per image and compute the standard deviation (std) of the resulting $K$ quality scores. As illustrated in Fig. 3, the std falls steadily across all datasets, indicating that predictions of VisualQuality-R1 become progressively more stable and confident.
@@ 5 Conclusion and Discussion
Despite the generalization capabilities demonstrated by VisualQuality-R1, several limitations and promising research directions merit further discussion. First, as a specific case of test-time scaling, VisualQuality-R1 is slow, expensive, and memory-hungry; it may also compound early errors into confidently wrong predictions. It is thus desirable to incorporate sample-adaptive reasoning, rationale compression or distillation, and self-consistency sampling to make VisualQuality-R1 faster, cheaper, and more robust. Second, VisualQuality-R1 relies on a single, fixed text prompt for all images, regardless of the underlying distortion scenario or application context. Incorporating application-aware prompt adaptation, for example via learned prompt-tuning or dynamic template selection, could tailor VisualQuality-R1’s reasoning and scoring to specific image processing tasks, therefore improving its flexibility and accuracy. Third, VisualQuality-R1 is currently formulated as an NR-IQA model, focusing solely on distorted inputs without access to pristine-quality counterparts. It is interesting to extend VisualQuality-R1 to a reference-based setting, which allows a (possibly corrupted) reference image wu2024comprehensive —potentially differing in resolution, color gamut, dynamic range, or bit depth—to serve as a flexible anchor for content fidelity. Last, we foresee adapting the proposed RL2R learning algorithm to other perceptual assessment tasks, including image aesthetics assessment tang2013content , human age estimation guo2009human , and perceptual similarity ranking fu2023dreamsim . Collectively, these promising directions aspire to foster more intelligent, transparent, and adaptable perceptual systems.
Acknowledgments
This work was supported in part by the Hong Kong ITC Innovation and Technology Fund (9440379 and 9440390) and the PolyU-OPPO Joint Innovative Research Center.
Appendix
Appendix A Generalization Probing via gMAD Competition
The group maximum differentiation (gMAD) competition ma2018group is a model comparison framework designed to evaluate the generalization capability of computational models, particularly in scenarios where exhaustive ground-truth labeling is impractical. In this framework, models take turns serving as the defender, while the remaining models act as attackers. The attackers aim to identify image pairs that the defender assigns similar quality scores to, but which they themselves rate very differently. These “adversarial” pairs are subsequently assessed by human observers, allowing for an efficient and targeted examination of model strengths and weaknesses.
Applying gMAD to compare KADID-10K-trained VisualQuality-R1 and Q-Insight on KonIQ-10k hosu2020koniq reveals a clear performance distinction. VisualQuality-R1 consistently uncovers perceptual inconsistencies in Q-Insight’s predictions. Conversely, when subjected to attacks, VisualQuality-R1 demonstrates strong robustness, aligning closely with human perception of image quality. This dual capability—high aggressiveness in exposing flaws and strong resistance to adversarial challenges—underscores VisualQuality-R1’s superior generalization, affirming its reliability for real-world IQA.
Appendix B Model Complexity Comparison
We present a quantitative comparison of model complexity across representative NR-IQA methods. As shown in Table 5, VisualQuality-R1 incurs higher inference costs than discriminative models such as MUSIQ ke2021musiq and MANIQA yang2022maniqa , as well as small-scale VLMs like LIQE zhang2023blind . This increased computational demand, however, reflects a deliberate design choice to support fine-grained quality reasoning and robust quality rating across diverse distortion scenarios.
Figure 5: gMAD competition results between VisualQuality-R1 and Q-Insight li2025qinsight . (a) Fixed Q-Insight at the low-quality level. (b) Fixed Q-Insight at the high-quality level. (c) Fixed VisualQuality-R1 at the low-quality level. (d) Fixed VisualQuality-R1 at the high-quality level.
Table 5: Model complexity comparison using a $512\times 384\times 3$ image as input.
Method #Parameters
Inference
Time
Inference
Memory
#TFLOPs
UNIQUE zhang2021uncertainty 22.32 M 0.02 s 1.62 G 0.029
MUSIQ ke2021musiq 27.13 M 0.05 s 1.69 G 0.026
MANIQA yang2022maniqa 135.75 M 0.03 s 2.13 G 0.217
LIQE zhang2023blind 151.28 M 0.03 s 2.15 G 0.131
Q-Align wu2024qalign 8.20 B 0.14 s 17.1 G 1.98
DeQA-Score you2025teaching 8.20 B 0.11 s 17.1 G 1.98
Q-Insight li2025qinsight 8.29 B 2.72 s 17.6 G 8.71
VisualQuality-R1 8.29 B 2.34 s 17.6 G 7.74
Figure 6: Visual reasoning samples from VisualQuality-R1 under the synthetic distortion scenario.
Figure 7: Visual reasoning samples from VisualQuality-R1 under the realistic distortion scenario.
Appendix C Additional Visual Reasoning Examples
We provide additional visual reasoning examples to illustrate the perceptual alignment of VisualQuality-R1 under both synthetic and realistic distortion scenarios. It should be noted that VisualQuality-R1 occasionally assigns quality scores outside the prescribed range of $\left[1,5\right]$ (e.g., a score of $0.2$ for the second example in Fig. 6), which reflects a common limitation in LLMs. Nevertheless, these deviations have negligible impact on the overall quality prediction performance.